OpenStudy (anonymous):

Find the cube of 20 A: 40 B: 400 C: 800 D: 8000

OpenStudy (calculusxy):

\[\large 20^3 = 20 \times 20 \times 20 \]

OpenStudy (anonymous):

ok thank you!

OpenStudy (calculusxy):

no problem!